OpenStudy (anonymous):

Look at the table of values below. x y 1 3 2 7 3 11 Which equation is represented by the table?

OpenStudy (anonymous):

y=2x+1 y=3x+2 y=4x-1 y=5x-3

OpenStudy (anonymous):

@jcpd910 @kirbykirby @phi

OpenStudy (kirbykirby):

You can find the slope of the line by finding the following: \[\frac{y_2-y_1}{x_2-x_1} \] where (\(x_1, y_1)\), \((x_2, y_2)\) are two coordinates given from your table. In fact, once you determine this, only one answer will be good since all the given slopes are different, so calculating the y-intercept would;t be necessary

OpenStudy (anonymous):

A little confusing

OpenStudy (kirbykirby):

But if you did want to find the y-intercept, to be certain, then you can just plug in any point from your table, and solve for \(b\)

OpenStudy (anonymous):

@jcpd910

OpenStudy (kirbykirby):

Say you choose two points (1, 3) and (2, 7) These will be \((x_1, y_1)\) and \((x_2, y_2)\).

OpenStudy (anonymous):

kirby explained it, well

OpenStudy (anonymous):

Im so confused

OpenStudy (anonymous):

7 - 3 over 2 - 1 4 over 1 That's your slope.

OpenStudy (anonymous):

So, 4 over 1 is 4

OpenStudy (anonymous):

and the equation is y = mx + b where m is the slope and b is the y intercept.

OpenStudy (anonymous):

if four is your slope, you can see that in your answer choices there is only one correct answer.

OpenStudy (anonymous):

Thank you very much @jcpd910. Thank you also @kirbykirby but you have to break it down a little better for me to understand this graphing stuff.

OpenStudy (anonymous):

No problem.